OpenStudy (anonymous):

Hi guys could i please get help on mechanical tehnology in the chapter on frictions......what is the meaning of the angle of friction?

OpenStudy (vincent-lyon.fr):

Angle of friction is the maximum angle between normal component and reaction force. tan φ = T/N where \(\vec R=\vec N +\vec T\)  In static friction T/N < tan φ = μstatic It happens to be the the maximum angle you can lift an inclined plane before an object starts slipping.